OpenStudy (mtalhahassan2):

Need Help!! Solve each equation using graphing technology. Round your answers to two decimal places, if necessary. a) 2x-5/x+10= 1/x-6

OpenStudy (anonymous):

not clear i sit \[\frac{2x-5}{x+10}=\frac{1}{x-6}\]?

OpenStudy (anonymous):

not so clear

OpenStudy (mtalhahassan2):

yep its right

OpenStudy (mtalhahassan2):